If the perimeter of a rectangle is 18 ft and the width is 6 ft, find its length. [Hint: The perimeter of a rectangle is given by P = 2L + 2W.]

Answer by Earlsdon(6294)

Starting with the formula for the perimeter of a rectangle:
P = 2L+2W and given that P = 18, W = 6, substitute these into the formula and solve for L
18 = 2L+2(6)
18 = 2L+12 Subtract 12 from both sides.
18-12 = 2L
6 = 2L Divide both sides by 2.
6/2 = 2L/2 Simplify.
3 = L or L = 3